OpenStudy (anonymous):

Prove that for each n in N, 5 divides n^5-n

OpenStudy (anonymous):

by induction, so if n=0 then 0/5?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so p(k) = k^5-k and p(k+1) = (k+1)^5-(k+1)

OpenStudy (anonymous):

I'm not sure how to manipulate it

ganeshie8 (ganeshie8):

familiar with binomial theorem ?

ganeshie8 (ganeshie8):

(k+1)^5 = k^5 + 5M + 1

ganeshie8 (ganeshie8):

you can see that if you expand (k+1)^5, all the middle terms will be multiples of 5

OpenStudy (ikram002p):

how this is helping ?

ganeshie8 (ganeshie8):

(k+1)^5-(k+1) = k^5 + 5M + 1 - (k+1) = k^5 - k + 5M

OpenStudy (ikram002p):

oh induction

ganeshie8 (ganeshie8):

from induction assumption k^5-k is divisible by 5, so the proof is over.

OpenStudy (anonymous):

so basically where it's +5M it shows that it is also divisible by 5?

ganeshie8 (ganeshie8):

yes, but i think it would make more sense if you expand it out

ganeshie8 (ganeshie8):

do you know how to expand (k+1)^5 ?

ganeshie8 (ganeshie8):

remember the (x+1)^n frormula or binomial theorem ?

OpenStudy (anonymous):

uh like, k^5+5k^4+10k^3+10k^2+5k-k

ganeshie8 (ganeshie8):

exactly, notice that all the middle terms have a factor 5 in common

OpenStudy (anonymous):

so they have to be divisible by 5! Perfect thanks!

ganeshie8 (ganeshie8):

(k+1)^5 = k^5 + 5k^4 + 10k^3 + 10k^2 + 5k + 1

ganeshie8 (ganeshie8):

you got it! but there are many other nice ways to prove this, are you supposed to do this using induction only is it ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

okay then your work looks good, ust make sure you justify every step

OpenStudy (anonymous):

Thank you :)

ganeshie8 (ganeshie8):

you're welcome!